A 60W-100V tungsten lamp has a resistance of 20 ohms at air temperature (0◦C). What is the rise in temperature of the filament under normal working conditions? The temperature coefficient of resistance of tungsten is 0.0052/◦C.

1
Expert's answer
2022-03-30T14:31:30-0400

The initial resistance

"R_0=20\:\Omega"

The resistance of the filament at the normal working condition

"R=\frac{V^2}{P}=\frac{100^2}{60}=167\:\Omega"

"R=R_0(1+\alpha t)"

"t=\frac{R-R_0}{\alpha}=\frac{167-20}{0.052}=2820\:\rm K"